40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16251     Entry time: Mon Jul 19 22:16:08 2021
Author: paco 
Type: Update 
Category: LSC 
Subject: PRFPMI locking 

[gautam, paco]

Gautam managed to lock PRFPMI a little before ~ 22:00 local time. The ALS to RF handoff logic was found to be repeatable, which enabled us to lock a total of 4 times this evening. Under this nominal state, we can work on PRFPMI to narrow down less known issues and carry out systematic optimization. The second time we achieved lock, we ran sensing lines before entering the ASC stage (which we knew would destroy the lock), and offline analysis of the sensing matrix is pending (gpstime = 1310792709 + 5 min).

Things to note:

(a) there is an unexpected offset suggesting that the ALS and RF disagreed on what the lock setpoint should be, and it is still unclear where the offset is coming from.

(b) the first time the lock was reached, the ASC up stage destroyed it, suggesting these loops need some care (we were able to engage the ASC loops at low gains (0.2 instead of 1) but as soon as we enabled some integrators this consistently destroyed the lock

(c) gautam had (burt) restored to the settings from back in March when the PRFPMI was last locked, suggesting there was a small but somehow significant difference in the IFO that helped today relative to last week


Take home message--> The mere fact that we were able to lock PRFPMI rules out the considerably more serious problems with the signal chain electronics or processing. This should also be a good starting point for further debugging and optimization.


gautam: the circulating power, when the ASC was tweaked, hit 400 (normalized to single arm locked with a misaligned PRM) suggesting a recycling gain of 22.5, and an average arm loss of ~30ppm round trip (assuming 2% loss in the PRC). 

ELOG V3.1.3-